### v1

# Why $w\neq-1$? Anthropic Selection in a $\Lambda$ + Axion Dark Energy Model

Abstract
We study a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers can exist and observe an accelerating universe, we derive nontrivial constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below because extremely light axions would require the bare cosmological constant to be unnaturally close to zero to achieve accelerated expansion. As a result, the anthropically allowed axion mass range typically lies around $m=\mathcal{O}(10)\,H_{0}$ for a decay constant close to the Planck scale, where $H_{0}$ is the observed value of the Hubble constant. In this framework, the dark energy equation of state parameter $w_{0}$ generically deviates from −$-1$ by $\mathcal{O}(0.1)$, providing a natural explanation for why $w\neq-1$ may be expected. This outcome is intriguingly consistent with recent DESI hints of time-varying dark energy, and offers a compelling anthropic explanation within the $\Lambda$ + axion framework.
I Introduction
The accelerated expansion of the universe is now well established through various cosmological observations, including Type Ia supernovae Perlmutter et al. (1999); Riess et al. (1998), the cosmic microwave background Aghanim et al. (2020), and large-scale structure Alam et al. (2017). This phenomenon is commonly attributed to dark energy, which constitutes roughly 70% of the current energy density of the universe. Despite its dominant role, the physical origin of dark energy remains one of the most profound mysteries in modern cosmology. A wide variety of theoretical models have been proposed to explain it, ranging from modifications of gravity Carroll et al. (2004) to dynamical scalar fields Ratra and Peebles (1988); Caldwell et al. (1998). Nevertheless, the simplest and most conservative candidate is a cosmological constant Weinberg (1989).
However, the cosmological constant faces a severe theoretical challenge. Naively, quantum field theory predicts a vacuum energy density of order $M_{\mathrm{Pl}}^{4}$, where $M_{\mathrm{Pl}}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass, while observations suggest a value smaller by about 120 orders of magnitude. This enormous discrepancy constitutes the so-called cosmological constant problem Weinberg (1989). In addition, string theory suggests that constructing metastable vacua with positive vacuum energy is nontrivial, while vacua with negative cosmological constant are far more generic Kachru et al. (2003); Bousso and Polchinski (2000); Obied et al. (2018); Danielsson and Van Riet (2018); Demirtas et al. (2022). Therefore, understanding the mechanism by which positive vacuum energy is uplifted from more generic negative-energy vacua remains a key challenge in connecting string theory to cosmological observations.
Among the proposed solutions to the cosmological constant problem, the anthropic principle Weinberg (1987, 1989) has gained significant attention, especially in the context of the string landscape Susskind (2003); Polchinski (2006); Ellis and Smolin (2009). If the landscape contains a vast number of vacua with different vacuum energies, it is natural to expect that observers can only exist in regions of the multiverse where the cosmological constant falls within a narrow anthropic window: too large a value would suppress galaxy formation, while a negative value would lead to rapid recollapse. This line of reasoning offers a statistical explanation for the observed smallness of the cosmological constant.111 In fact, the observed amount of dark energy is smaller than the anthropic upper bound, which may call for explanation Sorini et al. (2024). This is beyond the scope of this letter.
In this work, we go beyond the single-component picture in the anthropic argument and consider a scenario in which dark energy arises from the sum of a bare negative cosmological constant and the energy of an ultra-light axion. [^2] A related anthropic approach has been taken in the context of the strong CP problem, where the sum of a cosmological constant and an axion potential was considered as the effective cosmological constant Takahashi (2009); Kaloper (2019). Such axions are generically predicted in string theory, with masses spanning many orders of magnitude—a scenario often referred to as the axiverse Arvanitaki et al. (2010). If the axion is sufficiently light, it can remain displaced from its potential minimum over cosmological timescales and contribute to the present-day dark energy density.
Our goal is not to determine the full anthropically allowed parameter space of the $\Lambda$ + axion dark energy model, which would be highly nontrivial due to the many free parameters and the complexity of possible cosmological histories. Instead, we restrict our attention to scenarios resembling our universe, where intelligent observers can exist at some point in cosmic history and observe cosmic acceleration. Within this context, we investigate what can be said about the properties of the axion and the dark energy equation-of-state parameter.
In a universe where intelligent observers exist and observe an accelerating expansion, the total dark energy is required to lie below an anthropic upper bound on the cosmological constant. Although this condition does not directly constrain each component individually, we show that it nonetheless leads to nontrivial constraints on both the axion mass and the bare cosmological constant. Specifically, the axion mass is bounded from above because large masses would require the initial misalignment angle to be finely tuned near the hilltop in order to avoid early relaxation and recollapse. Conversely, the axion mass is bounded from below because, for too small masses, the bare cosmological constant must be tuned to be very close to zero to achieve accelerated expansion. These dual constraints result in a finite range of allowed axion masses and bare vacuum energies.
Remarkably, within this constrained parameter space, the dark energy equation of state parameter $w$ is generically predicted to deviate from −$-1$ by $\mathcal{O}(0.1)$. This deviation arises naturally from the anthropic selection and is consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b), providing a compelling explanation within the $\Lambda$ + axion framework. See Refs. Tada and Terada (2024); Yin (2024); Cortês and Liddle (2024); Bhattacharya et al. (2024a); Mukherjee and Sen (2024); Notari et al. (2024); Jia et al. (2025); Hernández-Almada et al. (2024); Bhattacharya et al. (2024b); Berbig (2025); Wolf et al. (2025); Chakraborty et al. (2025); Borghetto et al. (2025); Nakagawa et al. (2025); Lee et al. (2025); Colgáin et al. (2025); Santos et al. (2025) for various explanations of the DESI results.
II $\Lambda$ + Axion dark energy model
As a concrete realization of the anthropic scenario outlined in the introduction, we consider a model in which dark energy is composed of a negative cosmological constant, $\rho_{\Lambda}<0$, and a single axion field $\phi$. While the string axiverse generically contains multiple axions, we focus on a single axion field for simplicity. This approximation is justified if the axion mass spectrum is hierarchical: heavier axions would have already settled into their minima, while much lighter ones remain effectively frozen, and their energy densities can be effectively absorbed into $\rho_{\Lambda}$.
The axion potential is taken to be
$V(\phi)=m^{2}f^{2}\left[1+\cos\left(\frac{\phi}{f}\right)\right]\ ,$ (1)
where $m$ is the axion mass, $f$ is its decay constant, and we choose the origin, $\phi=0$, to be the maximum of the potential. We set $f=M_{\mathrm{Pl}}$ throughout, but our results remain largely unchanged as long as $f$ is somewhat below the Planck mass (e.g., $f\sim 10^{17}\,{\rm GeV}$). We will return to the case with smaller decay constants in a later discussion. We also assume that the axion field is spatially homogeneous.
For our analysis, we treat $\rho_{\Lambda}$, $\theta_{\mathrm{i}}\equiv\phi_{\mathrm{i}}/f$, and $m$ as independent random variables with flat priors for $\rho_{\Lambda}$ and $\theta_{\mathrm{i}}$, while fixing all other parameters to the observed values. Here, $\phi_{\mathrm{i}}$ is the initial value of $\phi$.
In an expanding universe, the axion follows the equation of motion
$\ddot{\phi}+3H\dot{\phi}-m^{2}f\sin\left(\frac{\phi}{f}\right)=0\ ,$ (2)
where the dots denote derivatives with respect to the cosmic time $t$, and $H$ is the Hubble parameter determined by the Friedmann equation,
$3M_{\mathrm{Pl}}^{2}H^{2}=\rho_{\mathrm{DE}}+\rho_{m}\equiv\rho_{\Lambda}+\rho_{\phi}+\rho_{m}\ .$ (3)
Here, $\rho_{m}$ and $\rho_{\phi}$ are the energy densities of the non-relativistic matter and the axion, respectively. We neglect the radiation component, which is subdominant during the epoch of interest.
To compare cosmic histories with different parameters, we define the present time $t_{0}$ by $\rho_{m}(t_{0})=\rho_{m,0}$, where $\rho_{m,0}=3M_{\mathrm{Pl}}^{2}H_{0}^{2}\Omega_{m}$ is the observed matter density at present. Here, $\Omega_{m}\simeq 0.3$ is the observed matter density parameter, and $H_{0}\simeq 67\,\mathrm{km/s/Mpc}$ is the observed Hubble constant. We normalize the scale factor as $a(t_{0})=1$ so that the matter density evolves as
$\rho_{m}=\rho_{m,0}\,a^{-3}\ .$ (4)
On the other hand, the axion energy density is given by
$\rho_{\phi}=\frac{1}{2}\dot{\phi}^{2}+V(\phi)\ .$ (5)
Anthropic considerations for a conventional cosmological constant impose an upper bound on its energy density, $\rho_{\Lambda}<\rho_{\mathrm{\Lambda}}^{\rm max}$, derived from the requirement that galaxies can form Weinberg (1987, 1989). In the following, we adopt the upper bound from the star formation history, $\rho_{\mathrm{\Lambda}}^{\rm max}=2306\rho_{\mathrm{DE,obs}}$ Sorini et al. (2024) with $\rho_{\mathrm{DE,obs}}=3M_{\mathrm{Pl}}^{2}H_{0}^{2}\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}\simeq 0.7$. A lower bound also follows from the requirement that the universe should not recollapse too soon after galaxy formation, allowing sufficient time for the development of intelligent life. In our scenario, the total dark energy density is given by the sum of a negative cosmological constant and the energy of an ultra-light axion, and it must lie within the following range:
$0<\rho_{\mathrm{DE,\mathrm{i}}}<\rho_{\mathrm{\Lambda}}^{\rm max},$ (6)
where $\rho_{\mathrm{DE,\mathrm{i}}}$ is the initial energy density of dark energy. We refer to this as a conditional anthropic bound. The lower bound reflects the need for the axion to uplift the negative vacuum energy and initially behave like a cosmological constant in order to drive accelerated expansion. The upper bound ensures that galaxy formation occurs before the axion starts oscillating. Furthermore, we impose the additional condition that the universe has not yet reached the turnaround point at the present time, to account for the time required after galaxy formation for intelligent life to emerge. This is a simplifying assumption, and our results remain essentially unchanged even if we adopt the less restrictive condition that the universe has not recollapsed by the present time or if we adopt a slightly different time $t_{0}$.
Since we fix $f=M_{\mathrm{Pl}}$, the Hubble parameter is always smaller than the axion mass $m$ during the dark-energy-dominated epoch. As a result, if $m$ is much larger than the Hubble parameter, the axion would have rolled down the potential by now, and the universe would have started to contract and recollapse unless the initial field value were sufficiently close to the hilltop. This degree of fine-tuning becomes increasingly severe for larger values of $|\rho_{\Lambda}|$, as a larger axion mass is required to uplift the total vacuum energy into the conditional anthropic window (6).
While we can evaluate the axion dynamics numerically, let us first obtain an approximate analytic description. We assume that the axion remains near the hilltop even at the present time. Then, $\rho_{\mathrm{DE}}$ is approximately given by
$\rho_{\mathrm{DE}}\simeq\rho_{\Lambda}+2m^{2}f^{2}\ .$ (7)
With this approximation, we can solve the Friedmann equation to obtain
$a(t)\simeq\left(\frac{\rho_{m,0}}{\rho_{\mathrm{DE}}}\right)^{1/3}\sinh^{2/3}\left(\frac{3}{2}H_{\mathrm{DE}}t\right)\ ,$ (8)
with
$H_{\mathrm{DE}}\equiv\sqrt{\frac{\rho_{\mathrm{DE}}}{3M_{\mathrm{Pl}}^{2}}}\ .$ (9)
Then, the equation of motion becomes
$\ddot{\phi}+3H_{\mathrm{DE}}\frac{\cosh\left(\frac{3}{2}H_{\mathrm{DE}}t\right)}{\sinh\left(\frac{3}{2}H_{\mathrm{DE}}t\right)}\dot{\phi}-m^{2}\phi\simeq 0\ ,$ (10)
where we used $\sin(\phi/f)\simeq\phi/f$ for $\phi\ll f$. With the initial condition
$\phi(t=0)=\phi_{\mathrm{i}}\ ,\quad\dot{\phi}(t=0)=0\ ,$ (11)
the equation of motion can be solved to yield
$\phi(t)$ $=\phi_{\mathrm{i}}\frac{\sinh(\sqrt{1+A}\tilde{t})}{\sqrt{1+A}\sinh(\tilde{t})}\ ,$ (12)
where
$A\equiv\frac{4m^{2}}{9H_{\mathrm{DE}}^{2}}\ ,\quad\tilde{t}\equiv\frac{3}{2}H_{\mathrm{DE}}t\ .$ (13)
Thus, at the present time, we obtain
$\phi(t_{0})=$ $\phi_{\mathrm{i}}\frac{\sinh(\sqrt{1+A}\tilde{t}_{0})}{\sqrt{1+A}\sinh(\tilde{t}_{0})}\ ,$ (14)
$\dot{\phi}(t_{0})=$ $\frac{3}{2}H_{\mathrm{DE}}\phi_{\mathrm{i}}$
×$\times\left[\frac{\cosh(\sqrt{1+A}\tilde{t}_{0})}{\sinh(\tilde{t}_{0})}-\frac{\cosh(\tilde{t}_{0})\cosh(\sqrt{1+A}\tilde{t}_{0})}{\sqrt{1+A}\sinh^{2}(\tilde{t}_{0})}\right]\ ,$ (15)
with
$\tilde{t}_{0}\equiv\frac{3}{2}H_{\mathrm{DE}}t_{0}=\mathrm{arcsinh}\left(\sqrt{\frac{\rho_{\mathrm{DE}}}{\rho_{m,0}}}\right)\ .$ (16)
The equation-of-state parameter at the present time $t=t_{0}$ is given by
$w_{0}$ $=\left.\frac{\frac{1}{2}\dot{\phi}^{2}-\rho_{\Lambda}-V(\phi)}{\frac{1}{2}\dot{\phi}^{2}+\rho_{\Lambda}+V(\phi)}\right|_{t=t_{0}}\ .$ (17)
We evaluate the evolution of the axion and $w_{0}$ both numerically and analytically. In Fig. 1, we show the parameter region allowed by the conditional anthropic bound for $m=100H_{0}$, $10H_{0}$, and $H_{0}$. We fix the range of $\rho_{\Lambda}$ following the condition (6). Note that the scales of the horizontal and vertical axes are different in the three panels. In the gray-shaded regions, the universe starts to contract before the present time, and we exclude these regions based on the requirement that the universe must not recollapse too soon after galaxy formation. We also show $w_{0}$ by the color shading for the allowed regions. Smaller values of $\theta_{\mathrm{i}}$ lead to smaller $w_{0}$, as the axion remains closer to the hilltop. The required tuning of $\theta_{\mathrm{i}}$ becomes more severe for small $\rho_{\Lambda}$ because the Hubble parameter is small, which allows the axion to roll down the potential more easily. Near the boundaries of the allowed region, $w_{0}$ becomes larger since the denominator of Eq. (17) becomes smaller. In the narrow regions at the edge of the allowed regions, the sign of $w_{0}$ flips due to the negative value of the denominator. While this is visible for $m=H_{0}$, such a region cannot be resolved for $m=100H_{0}$ due to the limited resolution of the plot.
Figure 1: Colored regions indicate the predicted values of $w_{0}$, and dashed lines show analytic estimates of the boundaries between them. The axion mass is taken to be $m=100H_{0}$, $10H_{0}$, and $H_{0}$ from top to bottom. The gray-shaded region in each panel is excluded because the universe recollapses before the present time.
We also show the results of the analytic estimate with dashed lines. Below the black dashed lines, $\rho_{m,0}+\rho_{\Lambda}+V(\phi(t_{0}))<0$ or $\phi(t_{0})>\pi f$ holds. Note that we excluded the kinetic energy from the total energy in the criterion for recollapse because it can be wrongly large for $\phi(t_{0})\gtrsim f$ in the analytic approximation. The colored dashed lines correspond to $w_{0}=-0.9,-0.6,-0.3,0$ basically from left to right. For $m=H_{0}$, there also appears an isolated region with $w<-0.9$ for the reason mentioned above. While the analytic estimate of $w_{0}$ agrees well with the numerical results for small $w_{0}$ and large $m$, it worsens in the opposite case, where the approximations of $\phi\ll f$ and $\rho_{\phi}=\mathrm{const}.$ break down.
With the flat priors for $\theta_{\mathrm{i}}$ and $\rho_{\Lambda}$, we can evaluate the likelihood of $\ln m$ by integrating the area of the colored regions in Fig. 1. We show the likelihood of $\ln m$ in Fig. 2. The vertical axis is normalized so that the maximum value equals unity. We find a peak at $m=m_{\mathrm{peak}}\simeq 49.2H_{0}$, which is determined by $2m^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$. This behavior can be understood as follows. For small $m$, the allowed value for $\rho_{\Lambda}$ is limited by −$-2m^{2}f^{2}<\rho_{\Lambda}<0$ by Eq. (6), and the likelihood increases roughly ∝$\propto m^{2}$. For large $m$, while the interval of the allowed $\rho_{\Lambda}$ becomes equal to $\rho_{\mathrm{\Lambda}}^{\mathrm{max}}$, independent of $m$, the required tuning of $\theta_{\mathrm{i}}$ near the hilltop becomes exponentially severe, suppressing the likelihood. While the peak position of the axion mass distribution depends on the values of $\rho_{\Lambda}^{\mathrm{max}}$ and $f$ as $2m_{\mathrm{peak}}^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$, the existence of the sharp peak itself does not depend on their values.
Figure 2: Likelihood of $\ln m$ marginalized over $\theta_{\mathrm{i}}$ and $\rho_{\Lambda}$. The vertical axis is normalized by the maximum value. Lower masses are disfavored due to the limited range of $\rho_{\Lambda}$, while higher masses are suppressed due to fine tuning of $\theta_{\mathrm{i}}$.
In Fig. 3, we show the probability distribution of $w_{0}$ for fixed $m$. Here, we sample $w_{0}$ from −$-1$ to $1$ in increments of $0.1$, and $w_{0}>1$ are collected into the rightmost bin. We evaluate the probability by integrating the area with corresponding values of $w_{0}$ in Fig. 1 and normalize it such that the total equals unity. We find that while the most likely value of $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant. Thus, our results suggest that, in a universe where intelligent observers can observe the accelerated expansion, the dark energy equation-of-state parameter is likely to deviate from −$-1$ at the level of ${\cal O}(0.1)$.
Figure 3: Histograms of $w_{0}$ for the three axion masses, $m=100H_{0}$, $10H_{0}$, and $H_{0}$, are shown in different colors.
III Discussion and Conclusions
So far, we have focused on the case where the bare cosmological constant is restricted to be negative. This simplifying assumption is well-motivated in the context of string theory, where vacua with negative vacuum energy are favored. Nonetheless, it is natural to consider generalizations in which the bare cosmological constant can also take positive values. We have also assumed that observers exist in the universe and observe cosmic acceleration. Relaxing this assumption is another possible extension. If one allows a bare positive cosmological constant or does not impose the condition that observers see an accelerating universe, the anthropic argument no longer provides a lower bound on $m$ when assuming a flat measure in $\ln m$ for the axion mass. This is because the probability distribution remains constant for arbitrarily small $m$, leading to a divergent probability. This requires some form of regularization, either through a modified measure or by imposing a prior on the axion mass.
It should also be noted that including both eternally expanding universes and those that eventually recollapse in the anthropic argument is itself nontrivial. In the present work, we imposed a lower bound on the recollapse time, assuming that the emergence of intelligent observers like ourselves requires a timescale after galaxy formation, comparable to the current age of our universe. Unlike the case of a positive cosmological constant, where galaxies can persist indefinitely once formed, a universe with negative vacuum energy inevitably undergoes recollapse within a finite time. Whether such recollapsing universes should be assigned the same anthropic weight as eternally expanding ones is unclear, and addressing this issue may require a more refined treatment of the multiverse measure. A proper assessment of these issues would involve a careful reevaluation of the prior distributions over both the cosmological constant and the axion parameters. We leave such investigations for future work.
In this Letter, we have explored a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers exist and observe an accelerating universe, a conditional anthropic bound (6) naturally follows, which constrains the initial total dark energy density $\rho_{\mathrm{DE}}$ to lie within a range compatible with both galaxy formation and cosmic acceleration. Under this condition, we derived nontrivial anthropic constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below to prevent the total dark energy from requiring an unnaturally small bare cosmological constant.
We have found that, within the anthropically allowed region, the axion mass typically lies around ${\cal O}(10)H_{0}$ for a decay constant near the Planck scale, and it increases for smaller decay constants. While the most likely value of the dark energy equation-of-state parameter $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant. This provides a natural anthropic explanation for why $w_{0}\neq-1$ may be expected, and is intriguingly consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b). Our results thus offer a concrete realization of how the string theory landscape, together with a light axion, can shape the nature of dark energy in our universe.
Acknowledgments
We thank Wen Yin for helpful discussion on the DESI results. This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T.), 23KJ0088 (K.M.), 24K17039 (K.M.), and 25H02165 (F.T.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).